# On the quantum chromatic number of Hamming and generalized Hadamard graphs

### v1

@@ 2 Preliminaries
All graphs considered in this paper are assumed to be simple, that is, undirected, with no loops or multiple edges. Directed graphs will be explicitly indicated when relevant.
Given two graphs $G$ and $H$, a homomorphism from $G$ to $H$ is a map $\phi:V(G)\to V(H)$ such that if $u$ and $v$ are adjacent in $G$, then $\phi(u)$ and $\phi(v)$ are adjacent in $H$. The chromatic number of $G$, denoted by $\chi(G)$, is the minimum positive integer $r$ such that there exists a homomorphism from $G$ to the complete graph on $r$ vertices. The clique number of $G$, denoted by $\omega(G)$, is the maximum positive integer $r$ such that there exists a homomorphism from the complete graph on $r$ vertices to $G$. The independence number of $G$, denoted by $\alpha(G)$, is the maximum size of an independent set in $G$ and is equal to the clique number of the complement of $G$. Observe that $\chi(G)\alpha(G)\geq|V(G)|$.
A quantum homomorphism from $G$ to $H$ is a set of Hermitian matrices
$\left\{P_{v,\alpha}:v\in V(G),\alpha\in V(H)\right\}\subseteq\mathbb{C}^{k\times k}$
for some positive integer $k$, satisfying:
@@
    For every $v\in V(G)$, the set $\{P_{v,\alpha}:\alpha\in V(H)\}$ forms a complete orthogonal system, which means that $\sum_{\alpha\in V(H)}P_{v,\alpha}=I$ and $P_{v,\alpha}P_{v,\beta}=\delta_{\alpha,\beta}P_{v,\alpha}$ for all $\alpha,\beta\in V(H)$, where $\delta_{\alpha,\beta}$ equals 1 if $\alpha=\beta$ and 0 otherwise.
@@
    For any two adjacent vertices $u,v\in V(G)$ and any two non-adjacent vertices $\alpha,\beta\in V(H)$, we have $P_{u,\alpha}P_{v,\beta}=O$.
Using quantum homomorphisms, the quantum chromatic number of $G$, denoted by $\chi_{Q}(G)$, is the minimum positive integer $r$ such that there exists a quantum homomorphism from $G$ to the complete graph on $r$ vertices. Clearly, any graph homomorphism naturally induces a quantum homomorphism. Therefore, we have $\chi_{Q}(G)\leq\chi(G)$.
We next recall two classical bounds for the quantum chromatic number: an upper bound based on modulus-one orthogonal representations and a spectral lower bound.
An orthogonal representation of a graph $G$ with dimension $K$ is a map $\rho:V(G)\to\mathbb{C}^{K}$ such that $\rho(u)$ and $\rho(v)$ are orthogonal with respect to the complex inner product for all adjacent vertices $u,v\in V(G)$. Moreover, the representation $\rho$ is called modulus-one if all coordinates of $\rho(v)$ have modulus one for every $v\in V(G)$. Let $\xi(G)$ and $\xi^{\prime}(G)$ denote the minimum dimension of an orthogonal representation and of a modulus-one orthogonal representation, respectively.
As established by Cameron et al. in [4], the quantum chromatic number of $G$ satisfies the following upper bound:
Lemma 2.1.
$\chi_{Q}(G)\leq\xi^{\prime}(G)$.
On the other hand, spectral techniques provide a lower bound for $\chi_{Q}(G)$. Let $G$ be a graph on $n$ vertices, and let $\lambda_{1}\geq\dots\geq\lambda_{n}$ denote the eigenvalues of its adjacency matrix. The following Hoffman-type bound was established in [8]:
Lemma 2.2.
$\chi_{Q}(G)\geq 1-\frac{\lambda_{1}}{\lambda_{n}}$.
It is well known that if $G$ is an $r$-regular graph, then $\lambda_{1}=r$ is the largest eigenvalue in absolute value. Note that both Hamming graphs and generalized Hadamard graphs are regular. More precisely, $H(n,q,d)$ is $(q-1)^{d}\binom{n}{d}-$regular, and $\varOmega_{n}^{(\mathbb{G})}$ is $\binom{n}{n/q,\dots,n/q}-$regular.
@@
The Cayley graph on a group $\mathbb{G}$ with a generating set $S\subseteq\mathbb{G}$, denoted by $\operatorname{Cay}(\mathbb{G},S)$, is the directed graph on vertex set $\mathbb{G}$ where two vertices $g$ and $h$ form a directed edge $(g,h)$ if and only if $g^{-1}h\in S$. In particular, if $S^{-1}=S$, then $\operatorname{Cay}(\mathbb{G},S)$ is an undirected graph.
@@
In fact, Hamming and composition graphs can be represented as Cayley graphs. Given a $q$-ary alphabet $\varSigma$, equip it with an arbitrary additive group structure $\mathbb{G}=(\varSigma,+)$. For any $n$-tuple $x\in\mathbb{G}^{n}$, define its Hamming weight $w_{\mathrm{H}}(x)$ as the number of coordinates not equal to $0_{\mathbb{G}}$, and its composition $\operatorname{comp}(x)=\boldsymbol{i}=(i_{g}:g\in\mathbb{G})$, where $i_{g}$ is the number of coordinates of $x$ equal to $g$. The Hamming graph $H(n,q,i)$ and the composition graph $H_{C}(n,\mathbb{G},\boldsymbol{i})$ can be represented as the Cayley graphs $\operatorname{Cay}(\mathbb{G}^{n},S_{i})$ and $\operatorname{Cay}(\mathbb{G}^{n},S_{\boldsymbol{i}})$, respectively, where $S_{i}=\{x\in\mathbb{G}^{n}:w_{\mathrm{H}}(x)=i\}$ and $S_{\boldsymbol{i}}=\{x\in\mathbb{G}^{n}:\operatorname{comp}(x)=\boldsymbol{i}\}$. In particular, when $q\mid n$, we simply write the composition $(n/q,\dots,n/q)$ in bold as $\boldsymbol{n/q}$, and the generalized Hadamard graph $\varOmega_{n}^{(\mathbb{G})}=H_{C}(n,\mathbb{G},\boldsymbol{n/q})$ can be represented as the Cayley graph $\operatorname{Cay}(\mathbb{G}^{n},S_{\boldsymbol{n/q}})$.
@@
Note that both Hamming and Composition graphs are Cayley graphs over an Abelian group. The eigenvalues of an Abelian Cayley graph can be expressed in an elegant way. Before presenting their eigenvalues, we first recall some basics of the characters of an Abelian group.
@@
It is well known that for $\mathbb{G}=\mathbb{Z}_{q}^{n}$, the character group is given by $\widehat{\mathbb{Z}_{q}^{n}}=\{\varphi_{x}:x\in\mathbb{Z}_{q}^{n}\}$, with $\varphi_{x}(y)=\zeta_{q}^{x\cdot y}$ for all $y\in\mathbb{Z}_{q}^{n}$. Similarly, for $\mathbb{G}=\mathbb{F}_{q}^{n}$, where $q$ is a power of a prime $p$, we have $\widehat{\mathbb{F}_{q}^{n}}=\{\psi_{\alpha}:\alpha\in\mathbb{F}_{q}^{n}\}$, and the characters are defined by $\psi_{\alpha}(\beta)=\zeta_{p}^{\operatorname{Tr}(\alpha\cdot\beta)}$ for all $\beta\in\mathbb{F}_{q}^{n}$, where $\operatorname{Tr}(\cdot)$ denote the trace function from $\mathbb{F}_{q}$ to $\mathbb{F}_{p}$. In both cases, the dot product $x\cdot y$ is defined as $\sum_{i=1}^{n}x_{i}y_{i}$, and $\zeta_{r}$ be the primitive $r$-th roots of unity.
@@
$K_{i}(j)=\sum_{k=0}^{i}(-1)^{k}(q-1)^{i-k}\binom{j}{k}\binom{n-j}{i-k}$.
This fact leads naturally to the following reciprocal property, which can be derived by a simple double counting argument:
@@
From the second expression of $\mu_{x}$, it is clear that it depends only on the composition of $x$. That is, if $\mathrm{comp}(x)=\boldsymbol{j}=(j_{g}:g\in\mathbb{G})$, then
@@
We define the generalized Krawtchouk polynomial with respect to $\mathbb{G}$  [^2] In the special case $\mathbb{G}=\mathbb{F}_{q}$, this polynomial coincides with the one studied by Sookoo [13]. as

### v2

@@ 2 Preliminaries
For any positive integer $t$, we denote $t]\coloneqq\{1,\dots,t\}$. Section 2.1 is intended to provide intuition regarding the quantum principles and the role of shared entanglement in the graph coloring game. Since our main results rely solely on the algebraic definition of quantum coloring (see Definition 2.2), this subsection may be omitted without loss of continuity.
2.1 Basic quantum mechanics
To formally define the quantum chromatic number, we first introduce the necessary notation from quantum mechanics. We restrict our attention to finite-dimensional Hilbert spaces and use the standard Dirac bra-ket notation.
Let $\mathbb{C}^{k}$ denote a column vector space and a column vector in $\mathbb{C}^{k}$ is denoted by a ket $|v\rangle$, and its conjugate transpose is denoted by a bra $\langle v|$. The inner product of two vectors $|v\rangle$ and $|w\rangle$ is defined by the product $\langle v||w\rangle$, simply denoted by $\langle v|w\rangle$. In quantum information, the state of a system (or simply a particle) is treated as a unit vector in $\mathbb{C}^{k}$ (i.e., a vector $|v\rangle$ satisfying $\langle v|v\rangle=1$). Such a system is usually referred to as a qubit when $k=2$.
Definition 2.1.
A set of matrices $\{P_{1},P_{2},\ldots,P_{t}\}$ acting on $\mathbb{C}^{k}$ is called a complete orthogonal system if:
  1. 1.
    each $P_{i}$ is idempotent, i.e., $P_{i}^{2}=P_{i}$;
  2. 2.
    the system is mutually orthogonal, i.e., $P_{i}P_{j}={0}$ for any $i\neq j$;
  3. 3.
    the system is complete, i.e., $\sum_{i\in[t]}P_{i}=I_{k}$.
Suppose $\mathbb{C}^{k}$ admits a direct sum decomposition $\mathbb{C}^{k}=V_{1}\oplus V_{2}\oplus\cdots\oplus V_{t}$. Let $P_{i}$ be the projective operator from $\mathbb{C}^{k}$ onto $V_{i}$ along this decomposition, defined by $P_{i}(v_{1}+\cdots+v_{t})=v_{i}$ for all $v_{j}\in V_{j}$ and $j\in[t]$. Then, the set $\{P_{i}:i\in[t]\}$ forms a complete orthogonal system. Conversely, given a complete orthogonal system acting on $\mathbb{C}^{k}$, we obtain a direct sum decomposition $\mathbb{C}^{k}=\mathrm{im}(P_{1})\oplus\cdots\oplus\mathrm{im}(P_{t})$. Specifically, an orthogonal decomposition of $\mathbb{C}^{k}$ corresponds to a complete orthogonal system of Hermitian projectors (in the context of quantum information, a Hermitian idempotent is usually called an orthogonal projector).
Assume an orthogonal decomposition $\mathbb{C}^{k}=V_{1}\perp\cdots\perp V_{t}$, corresponding to a complete set of orthogonal projectors $\{P_{i}:i\in[t]\}$. Consider a state $|v\rangle\in\mathbb{C}^{k}$ (a unit vector). The state $|v\rangle$ can be uniquely expressed as $|v\rangle=|v_{1}\rangle+\cdots+|v_{t}\rangle$ with $|v_{i}\rangle\in V_{i}$. When a measurement defined by this decomposition (or equivalently, by the system $\{P_{i}\}$) is performed, the state $|v\rangle$ collapses to the normalized vector $\frac{|v_{i}\rangle}{\||v_{i}\rangle\|}$ with probability $\||v_{i}\rangle\|^{2}$. Observing the outcome $i$ corresponds to observing that the state has collapsed into the subspace $V_{i}$. The probability distribution is valid since $\sum_{i\in[t]}\||v_{i}\rangle\|^{2}=\||v\rangle\|^{2}=1$. This process is known as a projective measurement.
Using the relation $|v_{i}\rangle=P_{i}|v\rangle$, we can express the probability and the post-measurement state in terms of the projectors $P_{i}$:
$\Pr(i)$ $=\||v_{i}\rangle\|^{2}=\langle v_{i}|v_{i}\rangle=\langle v|P_{i}^{\dagger}P_{i}|v\rangle=\langle v|P_{i}^{2}|v\rangle=\langle v|P_{i}|v\rangle,$
Post-state $=\frac{|v_{i}\rangle}{\||v_{i}\rangle\|}=\frac{P_{i}|v\rangle}{\sqrt{\langle v|P_{i}|v\rangle}}.$
For simplicity, we refer to the set $\{P_{i}:i\in[t]\}$ as a projective measurement. As background, projective measurements are sufficient to understand the quantum graph coloring, so we focus on them. (While the general framework of quantum operations encompasses POVMs and unitary evolutions, these lie beyond the scope of our current discussion.)
Suppose $k$ is a positive integer. For each $i\in[k]$, let $|i\rangle\in\mathbb{C}^{k}$ denote the column vector with a $1$ in the $i$-th position and 0 elsewhere. The orthonormal basis $\{|i\rangle:i\in[k]\}$ is referred to as the computational basis. The corresponding set of projectors $\{|i\rangle\langle i|:i\in[k]\}$ serves as a fundamental example of a projective measurement. Note that in a general projective measurement, the subspaces $V_{i}$ need not have the same dimension.
Consider performing two sequential measurements on a quantum state $|v\rangle\in\mathbb{C}^{k}$, defined first by the set $\{P_{i}:i\in[t]\}$ and subsequently by $\{Q_{j}:j\in[s]\}$. This procedure yields an outcome pair $(i,j)\in[t]\times[s]$. To determine the joint probability $\Pr(i,j)$, we analyze the process step by step. The probability of obtaining the first outcome $i$ is $\langle v|P_{i}|v\rangle$. Upon this outcome, the state $|v\rangle$ collapses to the normalized vector $|v^{\prime}\rangle=\frac{P_{i}|v\rangle}{\sqrt{\langle v|P_{i}|v\rangle}}$. According to the postulates of quantum mechanics, the conditional probability of obtaining the second outcome $j$ given $i$ is:
$\Pr(j|i)=\langle v^{\prime}|Q_{j}|v^{\prime}\rangle=\frac{1}{\langle v|P_{i}|v\rangle}\langle v|P_{i}^{\dagger}Q_{j}P_{i}|v\rangle=\frac{\langle v|P_{i}Q_{j}P_{i}|v\rangle}{\langle v|P_{i}|v\rangle}.$
Consequently, the joint probability is given by $\Pr(i,j)=\langle v|P_{i}Q_{j}P_{i}|v\rangle$. This formula holds generally: if $\langle v|P_{i}|v\rangle=0$, both the probability $\Pr(i)$ and the joint probability vanish (consistent with $P_{i}|v\rangle={0}$); otherwise, $\Pr(i,j)=\Pr(i)\Pr(j|i)$, which also yields the expression above.
If the two measurements commute (i.e., $P_{i}Q_{j}=Q_{j}P_{i}$ for all $i,j$), the expression simplifies to:
$\Pr(i,j)=\langle v|P_{i}Q_{j}|v\rangle.$
In this commuting case, the order in which the measurements $\{P_{i}\}$ and $\{Q_{j}\}$ are applied does not affect the joint probability distribution.
Two quantum systems can be combined to form a composite system. For instance, the tensor product of the spaces $\mathbb{C}^{k}$ and $\mathbb{C}^{\ell}$ yields the joint space $\mathbb{C}^{k}\otimes\mathbb{C}^{\ell}$. A bipartite state $|v\rangle\in\mathbb{C}^{k}\otimes\mathbb{C}^{\ell}$ describes a system of two particles that can be distributed to two spatially separated observers, Alice and Bob. Mathematically, this separation implies that Alice’s measurements are of the form $\{P_{i}\otimes I_{\ell}:i\in[t]\}$, while Bob’s are of the form $\{I_{k}\otimes Q_{j}:j\in[s]\}$. Since operators acting on disjoint subsystems commute, the sequential measurement principle derived above applies, implying that the order of measurement is irrelevant. The resulting joint probability is given by
$\Pr(i,j)=\langle v|P_{i}\otimes Q_{j}|v\rangle.$
For convenience, we denote Alice’s and Bob’s measurements as $\{P_{i}\}$ and $\{Q_{j}\}$, respectively.
2.2 Quantum coloring for graphs
We first formalize the notion of quantum coloring. As rigorously formulated by Cameron et al. [8], the quantum chromatic number $\chi_{Q}$ relates to the existence of a specific system of projectors.
Definition 2.2.
Let $G$ be a simple graph and $r$ be a positive integer. Consider a set of Hermitian matrices
$\left\{P_{v,\alpha}:v\in V(G),\alpha\in[r]\right\}\subseteq\mathbb{C}^{k\times k},$
where $r]=\{1,2,\ldots,r\}$ and $k$ is a positive integer. The set $\{P_{v,\alpha}:v\in V(G),\alpha\in[r]\}$ is called a quantum $r$-coloring of $G$ if it satisfies the following conditions:
@@ 2.2 Quantum coloring for graphs
    For every vertex $v\in V(G)$, the set $\{P_{v,\alpha}:\alpha\in[r]\}$ forms a complete orthogonal system (see Definition 2.1).
@@
    For any color $\alpha\in[r]$ and any two adjacent vertices $u,v\in V(G)$, we have the orthogonality condition $P_{u,\alpha}P_{v,\alpha}=O$.
A quantum $r$-coloring $\{P_{v,\alpha}\}_{v,\alpha}$ provides a perfect strategy for the graph coloring game using shared entanglement. Specifically, Alice and Bob share a maximally entangled state $|v\rangle\in\mathbb{C}^{k}\otimes\mathbb{C}^{k}$, defined as
$|v\rangle=\frac{1}{\sqrt{k}}\sum_{i=1}^{k}|i\rangle\otimes|i\rangle.$
The two particles of $|v\rangle$ are distributed to Alice and Bob, respectively. Upon receiving vertices $x$ and $y$ respectively, Alice performs the measurement $\{P_{x,i}\}_{i}$ and Bob performs the measurement $\{\overline{P_{y,j}}\}_{j}$ (where $\overline{P}$ denotes the complex conjugate) on their own particles. Alice then reports her measurement result $\alpha$ to the verifier, and Bob similarly reports his measurement result $\beta$.
The probability that Alice reports color $\alpha$ and Bob reports color $\beta$ is given by
$\Pr(\alpha,\beta)$ $=\langle v|(P_{x,\alpha}\otimes\overline{P_{y,\beta}})|v\rangle$
$=\frac{1}{k}\left(\sum_{i\in[k]}\langle i|\otimes\langle i|\right)(P_{x,\alpha}\otimes\overline{P_{y,\beta}})\left(\sum_{j\in[k]}|j\rangle\otimes|j\rangle\right)$
$=\frac{1}{k}\sum_{i,j\in[k]}(\langle i|P_{x,\alpha}|j\rangle)(\langle i|\overline{P_{y,\beta}}|j\rangle)$
$=\frac{1}{k}\sum_{i,j\in[k]}\langle i|P_{x,\alpha}|j\rangle\langle j|P_{y,\beta}|i\rangle\quad\text{(since $P_{y,\beta}$ is Hermitian)}$
$=\frac{1}{k}\sum_{i\in[k]}\langle i|P_{x,\alpha}\left(\sum_{j\in[k]}|j\rangle\langle j|\right)P_{y,\beta}|i\rangle$
$=\frac{1}{k}\sum_{i\in[k]}\langle i|P_{x,\alpha}P_{y,\beta}|i\rangle=\frac{1}{k}\mathrm{Tr}(P_{x,\alpha}P_{y,\beta}).$
The conditions in Definition 2.2 ensure that this strategy wins with probability $1$. First, if the vertices $x$ and $y$ are adjacent ($x\sim y$), the probability of obtaining the same color $\alpha=\beta$ is
$\Pr(\alpha,\alpha)=\frac{1}{k}\mathrm{Tr}(P_{x,\alpha}P_{y,\alpha})=0,$
which follows from the orthogonality condition $P_{x,\alpha}P_{y,\alpha}=O$. Second, if the vertices are the same ($x=y$), the probability of obtaining different colors $\alpha\neq\beta$ is
$\Pr(\alpha,\beta)=\frac{1}{k}\mathrm{Tr}(P_{x,\alpha}P_{x,\beta})=0,$
since $\{P_{x,\gamma}\}_{\gamma}$ forms an orthogonal system.
Cameron et al. [8] showed that a quantum protocol employing general measurements and shared entanglement used to win the graph coloring game turns out to be a strategy utilizing projective measurements and maximal entangled state as described above. Building on this equivalence, the quantum chromatic number of a graph $G$, denoted by $\chi_{Q}(G)$, can be defined as the minimum integer $r$ for which a quantum $r$-coloring of $G$ exists.
For the reader’s convenience, we restate the following lemmas, which link the quantum chromatic number to the orthogonal rank and the spectrum of graphs, respectively.
Lemma 1.1 ([8, Proposition 7]).
For any graph $G$, $\chi_{Q}(G)\leq\xi^{\prime}(G)$.
Proof.
Let $k=\xi^{\prime}(G)$. Suppose $\rho:V(G)\rightarrow\mathbb{C}^{k}$ is a modulus-one orthogonal representation of $G$.
Denote $\zeta_{k}=e^{\frac{2\pi\sqrt{-1}}{k}}$. For each $x\in V(G)$ and $\alpha\in[k]$, define $P_{x,\alpha}=\Lambda_{x}|\alpha\rangle\langle\alpha|\Lambda_{x}^{-1}$, where $\Lambda_{x}=\operatorname{diag}(\rho(x)_{i}:i\in[k])$ and
$|\alpha\rangle=\frac{1}{\sqrt{k}}\begin{bmatrix}\zeta_{k}^{\alpha 0}\\ \zeta_{k}^{\alpha 1}\\ \vdots\\ \zeta_{k}^{\alpha(k-1)}\end{bmatrix}$.
Clearly, each $P_{x,\alpha}$ is Hermitian and $\{P_{x,\alpha}:\alpha\in[k]\}$ forms a complete orthogonal system for each $x\in V(G)$. It remains to verify the orthogonality condition, namely, $P_{x,\alpha}P_{y,\alpha}=0$ for all adjacent $x,y\in V(G)$, and $\alpha\in[k]$. We have
$P_{x,\alpha}P_{y,\alpha}=(\Lambda_{x}|\alpha\rangle\langle\alpha|\Lambda_{x}^{-1})(\Lambda_{y}|\alpha\rangle\langle\alpha|\Lambda_{y}^{-1})=0,$
since
$\langle\alpha|\Lambda_{x}^{\dagger}\Lambda_{y}|\alpha\rangle=\begin{bmatrix}\dots\ \zeta_{k}^{-\alpha i}\ \dots\end{bmatrix}\begin{bmatrix}\ddots&&\\ &\rho(x)_{i}^{*}&\\ &&\ddots\end{bmatrix}\begin{bmatrix}\ddots&&\\ &\rho(y)_{i}&\\ &&\ddots\end{bmatrix}\begin{bmatrix}\vdots\\ \zeta_{k}^{\alpha i}\\ \vdots\end{bmatrix}=\sum_{i=0}^{k-1}\rho(x)_{i}^{*}\rho(y)_{i}=0.$
Therefore, $\{P_{x,\alpha}:x\in V(G),\alpha\in[k]\}$ is a quantum $k$-coloring of $G$, hence $\chi_{Q}(G)\leq k=\xi^{\prime}(G)$. ∎
Lemma 1.2 ([16]).
Let $G$ be a graph with at least one edge. Then, $\chi_{Q}(G)\geq 1+\frac{\lambda_{1}}{|\lambda_{n}|}$.
2.3 Krawtchouk polynomials
We first recall some standard graph parameters. Given two graphs $G$ and $H$, a homomorphism from $G$ to $H$ is a map $\phi:V(G)\to V(H)$ such that if $u$ and $v$ are adjacent in $G$, then $\phi(u)$ and $\phi(v)$ are adjacent in $H$. The chromatic number of $G$, denoted by $\chi(G)$, is the minimum positive integer $r$ such that there exists a homomorphism from $G$ to the complete graph on $r$ vertices. The clique number of $G$, denoted by $\varOmega(G)$, is the maximum positive integer $r$ such that there exists a homomorphism from the complete graph on $r$ vertices to $G$. The independence number of $G$, denoted by $\alpha(G)$, is the maximum size of an independent set in $G$ and is equal to the clique number of the complement of $G$. It is noteworthy that $\chi(G)\alpha(G)\geq|V(G)|$.
@@ 2.3 Krawtchouk polynomials
Let $\mathbb{G}$ be a group and $S\subseteq\mathbb{G}$ a subset. The Cayley graph $\operatorname{Cay}(\mathbb{G},S)$ is the directed graph with vertex set $\mathbb{G}$, where a directed edge $(g,h)$ exists if and only if $g^{-1}h\in S$. If $S$ is inverse-closed (i.e., $S^{-1}=S$), then $\operatorname{Cay}(\mathbb{G},S)$ is undirected.
@@
Both Hamming and composition graphs can be realized as Cayley graphs over abelian groups. Let $\varSigma$ be a $q$-ary alphabet endowed with an additive group structure $\mathbb{G}=(\varSigma,+)$. For a tuple $x\in\mathbb{G}^{n}$, the Hamming weight $w_{\mathrm{H}}(x)$ is the number of coordinates not equal to $0_{\mathbb{G}}$, and the composition $\operatorname{comp}(x)$ is a tuple $\boldsymbol{i}=(i_{g}:g\in\mathbb{G})$ where $i_{g}$ is the number of coordinates of $x$ equal to $g$. The Hamming graph $H(n,q,d)$ and the composition graph $H_{C}(n,\mathbb{G},\boldsymbol{i})$ are isomorphic to the Cayley graphs $\operatorname{Cay}(\mathbb{G}^{n},S_{d})$ and $\operatorname{Cay}(\mathbb{G}^{n},S_{\boldsymbol{i}})$, respectively, where
$S_{d}=\{x\in\mathbb{G}^{n}:w_{\mathrm{H}}(x)=d\}\text{~~and~~}S_{\boldsymbol{i}}=\{x\in\mathbb{G}^{n}:\operatorname{comp}(x)=\boldsymbol{i}\}.$
In particular, when $q\mid n$, we denote the balanced composition $(n/q,\dots,n/q)$ by $\boldsymbol{n/q}$. Note that the generalized Hadamard graph over $\mathbb{G}$ (with $|\mathbb{G}|=q$) of length $n$ is defined as $\varOmega_{n}^{(\mathbb{G})}=H_{C}(n,\mathbb{G},\boldsymbol{n/q})$, which is therefore isomorphic to $\operatorname{Cay}(\mathbb{G}^{n},S_{\boldsymbol{n/q}})$.
@@
Recall that both Hamming and Composition graphs are Cayley graphs over an abelian group. The eigenvalues of an abelian Cayley graph can be expressed elegantly. Before presenting their eigenvalues, we first recall some basics of the characters of an abelian group.
@@
It is well known that for $\mathbb{G}=\mathbb{Z}_{q}^{n}$, the character group is given by $\widehat{\mathbb{Z}_{q}^{n}}=\{\varphi_{x}:x\in\mathbb{Z}_{q}^{n}\}$, with $\varphi_{x}(y)=\zeta_{q}^{x\cdot y}$ for all $y\in\mathbb{Z}_{q}^{n}$. Similarly, for $\mathbb{G}=\mathbb{F}_{q}^{n}$, where $q$ is a power of a prime $p$, we have $\widehat{\mathbb{F}_{q}^{n}}=\{\psi_{\alpha}:\alpha\in\mathbb{F}_{q}^{n}\}$, and the characters are defined by $\psi_{\alpha}(\beta)=\zeta_{p}^{\operatorname{Tr}(\alpha\cdot\beta)}$ for all $\beta\in\mathbb{F}_{q}^{n}$, where $\operatorname{Tr}(\cdot)$ denote the trace function from $\mathbb{F}_{q}$ to $\mathbb{F}_{p}$. In both cases, the dot product $x\cdot y$ is defined as $\sum_{i=1}^{n}x_{i}y_{i}$, and $\zeta_{r}$ is a primitive $r$-th root of unity.
@@
$K_{i}(j)=\sum_{k=0}^{i}(-1)^{k}(q-1)^{i-k}\binom{j}{k}\binom{n-j}{i-k}$. (1)
This fact leads naturally to the following reciprocal property, which a simple double-counting argument can derive:
@@
Another expression for $\mu_{x}$ shows that its value is determined solely by the composition of $x$. Specifically, if $\mathrm{comp}(x)=\boldsymbol{j}=(j_{g}:g\in\mathbb{G})$, then
@@
We define the generalized Krawtchouk polynomial over $\mathbb{G}$  [^1] In the special case $\mathbb{G}=\mathbb{F}_{q}$, this polynomial coincides with the one studied by Sookoo [29]. as